OpenStudy (taism8):

Trig/precal help. have to leave soon, need help asap <3

OpenStudy (taism8):

Find sin 2x, cos 2x, and tan 2x from the given information. sin x = \[\frac{ 8 }{ 17 }\] and x in Quadrant I sin 2x = cos 2x = tan 2x =

OpenStudy (taism8):



OpenStudy (taism8):

sinx = \[\frac{ 8 }{ 17 }\] cos x = \[\frac{ 15 }{17 }\] tan x = \[\frac{ 8 }{15 }\]

OpenStudy (taism8):

this is how much I know, I dont know how to do the sin2x cos2x and tan2x lol

OpenStudy (taism8):

@zepdrix

OpenStudy (taism8):

@Nnesha

zepdrix (zepdrix):

You figured out that the missing side length is 15 or something? Ok great. That's pretty much all we need.

OpenStudy (taism8):

ok

zepdrix (zepdrix):

From here, apply your Double angle formulas.\[\large\rm \sin2x=2\sin x \cos x\]And then plug those ratios in separately.

OpenStudy (taism8):

so it would be 2 * 8/17 * 15/17

zepdrix (zepdrix):

Mmmm yes good.

OpenStudy (taism8):

okie dokie

zepdrix (zepdrix):

Remember your other double angle identities? :)

OpenStudy (taism8):

are you sure tho? that doesnt seem right :/

OpenStudy (taism8):

yea google is my friend lol

OpenStudy (taism8):

simplifies to 612/289 which im not sure if my teacher would make such an absurd answer